# Polyp SAM 2: Advancing Zero shot Polyp Segmentation in Colorectal Cancer Detection

### v1

@@ Abstract
Polyp segmentation plays a crucial role in the early detection and diagnosis of colorectal cancer. However, obtaining accurate segmentations often requires labor-intensive annotations and specialized models. Recently, Meta AI Research released a general Segment Anything Model 2 (SAM 2), which has demonstrated promising performance in several segmentation tasks. In this work, we evaluate the performance of SAM 2 in segmenting polyps under various prompted settings. We hope this report will provide insights to advance the field of polyp segmentation and promote more interesting work in the future. This project is publicly available at https://github.com/sajjad-sh33/Polyp-SAM-2.
@@ 1 Introduction
The Segment Anything Model (SAM) [18] has demonstrated remarkable success in zero-shot image segmentation tasks. SAM’s prompt-based approach allows users to specify objects of interest without the need for additional training. By producing high-quality object masks from input prompts (such as points, boxes, and masks), SAM has garnered attention for various applications. For example, [26] assessed SAM’s zero-shot capabilities in organ segmentation tasks. [31] specifically evaluated SAM’s performance in Camouflaged Object Detection (COD). Meanwhile, [16] explored the benefits and limitations of the Segment-Anything Model (SAM) in both computer vision and medical image segmentation tasks. Additionally, [46] evaluated the performance of SAM in segmenting polyps, in which SAM is under unprompted settings. Deng et al. [5] applied SAM to segment heterogeneous objects in digital pathology.
The Segment Anything Model 2 (SAM 2) [7] builds upon the success of its predecessor, SAM, addressing the challenges posed by real time image and video segmentations. Trained on video data, SAM 2 offers real-time capabilities, allowing it to segment entire video sequences based on annotations from a single frame. By leveraging interactions (such as clicks, boxes, or masks) across frames, SAM 2 predicts spatial-temporal masks (referred to as ‘masklets’) for objects. Its unified approach reduces user interaction time and enhances performance, making it a promising tool for various applications, including medical image segmentation and video analysis [36, 27, 49].

### v2

@@ Abstract
Polyp segmentation plays a crucial role in the early detection and diagnosis of colorectal cancer. However, obtaining accurate segmentation often requires labour-intensive annotations and specialized models. Recently, Meta AI Research released a general Segment Anything Model 2 (SAM 2), which has demonstrated promising performance in several segmentation tasks. In this manuscript, we evaluate the performance of SAM 2 in segmenting polyps under various prompted settings. We hope this report will provide insights to advance the field of polyp segmentation and promote more interesting work in the future. This project is publicly available at https://github.com/sajjad-sh33/Polyp-SAM-2.
@@ 1 Introduction
The Segment Anything Model (SAM) [17] has demonstrated remarkable success in zero-shot image segmentation tasks. SAM’s prompt-based approach allows users to specify objects of interest without the need for additional training. By producing high-quality object masks from input prompts (such as points, boxes, and masks), SAM has garnered attention for various applications. For example, [26] assessed SAM’s zero-shot capabilities in organ segmentation tasks. Reference [31] specifically evaluated SAM’s performance in Camouflaged Object Detection (COD). Meanwhile, Reference [36] explored the benefits and limitations of the SAM in both computer vision and medical image segmentation tasks. Additionally, [46] evaluated the performance of SAM in segmenting polyps, in which SAM is under unprompted settings. Deng et al. [5] applied SAM to segment heterogeneous objects in digital pathology.
The Segment Anything Model 2 (SAM 2) [7] builds upon the success of its predecessor, SAM, addressing the challenges posed by real time image and video segmentation. Trained on video data, SAM 2 offers real-time capabilities, allowing it to segment entire video sequences based on annotations from a single frame. By leveraging interactions (such as clicks, boxes, or masks) across frames, SAM 2 predicts spatial-temporal masks (referred to as ‘masklets’) for objects. Its unified approach reduces user interaction time and enhances performance, making it a promising tool for various applications, including medical image segmentation and video analysis [36, 27, 49, 23].